# Evaluating Multi-Hop Reasoning in Large Language Models: A Chemistry-Centric Case Study

### v1

@@ 5.1 Context and Reasoning
In this analysis, we compare the performance of reasoning and non-reasoning models—i.e., models with and without test-time reasoning capabilities, respectively—in two scenarios: with context provided and without context provided. As illustrated in Figure 5-A, providing context significantly improves model performance, nearly doubling the correctness of both reasoning and non-reasoning models. Additionally, reasoning models outperform non-reasoning models in correctly answering questions, benefiting further from their reasoning capabilities in the context-provided setup. As expected and shown in Figure 5B, non-reasoning models are faster than reasoning models. Although these models benefit significantly from context to improve their performance, their latency did not significantly change by providing context to the model. This could be explained by the fact that the number of output tokens in these models is not sensitive to the length of the input prompt and availability of context (see Figure S6). For reasoning models, we observed that the availability of context lowers latency and output token count, likely because available context streamlines the thought process. For further analysis of context and reasoning in these models, refer to Appendix Evaluating Multi-Hop Reasoning in Large Language Models: A Chemistry-Centric Case Study.
@@
Figure 6: Analysis of the impact of the number of hops on models performance in Context Provided setup. panel A illustrates the relationship between output token usage (x-axis) and correctness rate (y-axis) for varying numbers of hops. Each point represents a model, while the colored areas indicate distribution estimates for each hop count. Panel B depicts the distribution of correctness rates for non-reasoning models across questions with different numbers of hops, with the dots representing the median of each distribution.
@@ 5.2 Impact of the Number of Hops
In this section, we investigate how the number of reasoning hops influences the correctness rate and the output token count. Figure 6 presents the results for the setup with Context Provided. For reasoning models, Figure 6-A illustrates the distribution of answer correctness in relation to the number of generated output tokens. The first observation is that as the number of hops increases, the output token count, which reflects the number of thinking tokens, also increases. However, the answer correctness rate remains relatively constant for multi-hop questions, albeit slightly lower than that observed in single-hop scenarios. For single-hop questions with context provided, we see a decrease in the correctness rate as the output token count increases, indicating a performance trade-off associated with deeper reasoning in simple questions. These trends are not present when context is not provided to the models, as shown in Supplementary Figure S7-A). In non-reasoning models, the output token count is not sensitive to the context or complexity of the question; thus, these models are evaluated solely based on answer correctness. Figure 6-B depicts the distribution of answer correctness rate across the evaluated reasoning models for different numbers of hops. Single-hop questions are answered at a higher correctness rate than multi-hop questions, a trend that is not observed in the setup without context (Supplementary figure S7-B).
@@ 6 Conclusion
In this study, we developed a domain-specific multi-hop question-answering (QA) system and evaluated state-of-the-art large language models within the chemistry domain. Our findings reveal that these models struggle with in-domain multi-hop scientific questions, correctly answering fewer than half of the queries when the context is unavailable. Although reasoning fine-tuned models show marginally improved performance, they still face significant challenges. The incorporation of context leads to substantial enhancements, nearly doubling the performance of both reasoning and non-reasoning models. However, even with context, no model, including those fine-tuned for reasoning, achieved a perfect score. Additionally, we contribute to the field by proposing an automated pipeline that integrates advanced named entity recognition with knowledge graph construction to generate intricate multi-hop reasoning tasks, which were utilized for the benchmark. Notably, this potentially domain-agnostic framework can be adapted for various fields by replacing chemistry-specific named entity recognition with suitable alternatives, laying a robust foundation for future research focused on improving reasoning capabilities across diverse specialized domains.
@@
Like any research, our study has certain limitations. Specifically, our benchmarking was conducted in two setups: without context and with full context provided. However, in real-world applications, each piece of context is typically collected step by step throughout the reasoning process. This raises the possibility of partial context retrieval, especially for non-reasoning models, which generally rely on a single retrieval round prior to generating answers. Equipping language models with a retrieval-augmented generation pipeline and allowing for multi-step retrieval could introduce a third scenario that better resembles real-world applications. In the past couple of years, advancements in reasoning models have led to various approaches for integrating generative models with RAG systems for multi-step retrieval [28, 37, 38, 39].
Building on these developments, future research could focus on designing and validating an industrial-grade, multi-step retrieval pipeline tailored to chemical text. This enhanced system will support incremental context acquisition and structured reasoning over lengthy or loosely organized documents, thereby enabling accurate answers to more challenging chemical queries.
@@ Acknowledgements
The author(s) gratefully acknowledge the financial support provided by MITACS under funding number IT32409 for the research leading to the publication of this article. We also thank Adam Wojciech Bartwiki for project management support, Tobias Roth for his essential contributions to the infrastructure, and Stephen Dokas for his invaluable recommendations in the chemistry domain.
Appendix
Appendix S1 Detailed Performance Based on Context Availability
@@
Figure S1 shows that model performance is strongly influenced by whether context is provided in the input. In particular, Claude 3.7 with extended thinking achieved a correctness rate of 84% with context, whereas o3-mini recorded the highest correctness rate (48%) when the context was absent. Note that o3-mini was primarily used to generate the questions, which may have introduced a slight bias resulting in its minor improvement in correctness. Figures S2 and S3 illustrate the token usage and latency of the models, respectively.
@@
S1.1 Reasoning without Provided Context
Figure S4 illustrates the performance of various models when no context is provided. As shown, o3-mini achieves the highest correctness rate, although this comes at a higher cost. It is important to note that o3-mini was primarily used for question generation, which may provide it with a slight advantage. In contrast, Llama3-3 70B Instruct records the lowest cost and lowest correctness rate while also exhibiting low latency. Meanwhile, Claude Sonnet 3.7 (without thinking tokens) balances correctness and cost, with acceptably low latency. This model also exhibited the highest correctness rate among non-reasoning models.
Figure S4: Performance of selected models across correctness rate, cost, and latency. The cost axis is log-scaled to highlight differences. The y-axis represents the average performance of models in providing the correct answer, and the dot size indicates model latency.
Appendix S2 Performance of models on Chemistry Subset of HotpotQA
Table S1 shows the details of each model performance, latency, and tokens used in both setups of context provided and not provided for the chemistry subset of HotpotQA [19] questions.
@@
Table S1: Summary of tested models performance on HotpotQA chemistry subset in terms of several evaluation metrics for both Contextual and Non-Contextual Setups
Appendix S3 A Multi-Hop QA Generation Example
@@
Carbonylation reactions constitute a potent tool to manufacture carboxylic acids and their derivatives both in industry and academic organic synthesis. In general, carbonylation requires the use of toxic carbon monoxide, which thus usually demands certified high-pressure reaction vessels. Therefore, developing non-gaseous CO surrogate for conducting safe and facile-operation carbonylation is an important and ongoing research topic. Among these established CO surrogates, formic acid is one kind of versatile atom.
The utilization of carbon dioxide as a C1 feedstock for the generation of industrially relevant chemicals is also an interesting approach. CO2 is an attractive renewable C1 source, which can lead to formic acid. Those approaches would not only reduce carbon dioxide emissions through carbon capture but also compensate sequestration costs by producing chemicals in global demand. Question:
@@
1) formic acid is the simplest carboxylic acid with antibacterial and preservative properties.
2) formic acid can be produced from carbon dioxide. 3) formic acid can act as a non-gaseous CO surrogate. 4) carbonylation reactions safely produce carboxylic acids under mild conditions using formic acid as a CO surrogate. Path (multi-hop chain of reasoning):
@@
Appendix S4 Impact of Context and Reasoning on output tokens count
Figure S6 illustrates the impact of context availability on the average number of output tokens generated by reasoning and non-reasoning models when answering questions. Non-reasoning models produce a similar number of tokens, as they do not engage in test-time reasoning. In contrast, for reasoning models, the number of tokens generated to answer questions decreases with the availability of context, indicating a potential requirement of less thinking when the context is available.
@@
S4.1 Performance Analysis Based on Number of Hops
Figure S7 illustrates how the number of hops affects performance in the absence of context. The data reveals that as the number of hops increases, the correctness rate declines while the number of output tokens rises. Additionally, both Figure LABEL:fig:hops_context and Figure S7 show a negative correlation between token count and correctness when only one hop is used. This may suggest that overanalyzing simpler questions could lead to errors in the final answer. We visualized the overall performance of all models in Figure S8 and  S9 to analyze how the correctness rate varies with the number of reasoning hops. In addition, token usage and latency metrics were separately depicted in Figures S10,  S11,  S12 and S13, respectively, to provide a more detailed view of the efficiency and resource consumption as the reasoning depth increases.
Figure S7: Analysis of the impact of the number of hops on model performance in the absence of context. A) Scatter and KDE plots showing the relationship between the number of hops, correctness rate, and output tokens. B) Distribution of non-reasoning models’ performance across different numbers of hops.
The following figures illustrate each model’s performance, latency, and token usage for question clusters that require different numbers of hops to answer.

### v2

@@ 5.2 Context and Reasoning
In this analysis, we compare the performance of reasoning and non-reasoning models, i.e., models with and without test-time reasoning capabilities, respectively, in two scenarios: with context provided and without context provided. As illustrated in Figure 5-A, providing context significantly improves model performance, nearly doubling the correctness of both reasoning and non-reasoning models. Additionally, reasoning models outperform non-reasoning models in correctly answering questions, benefiting further from their reasoning capabilities in the context-provided setup. As expected and shown in Figure 5B, non-reasoning models are faster than reasoning models. Although these models benefit significantly from context to improve their performance, their latency did not significantly change when context was provided to the model. This could be explained by the fact that the number of output tokens in these models is not sensitive to the length of input prompt and availability of context (see Figure S12).
For reasoning models, we observed that the availability of context lowers latency and output token count, likely because available context streamlines the thought process. For further analysis of context and reasoning in these models, refer to Appendix 7.
@@ 5.3 Impact of the Number of Hops
In this section, we investigate how the number of reasoning hops influences the correctness rate and the output token count. Figure 6 presents the results for the setup with Context Provided. For reasoning models, Figure 6-A illustrates the distribution of answer correctness in relation to the number of generated output tokens. The first observation is that as the number of hops increases, the output token count, which reflects the number of thinking tokens, also increases. However, the answer correctness rate remains relatively constant for multi-hop questions, albeit slightly lower than that observed in single-hop scenarios. For single-hop questions with context provided, we see a decrease in the correctness rate as the output token count increases, indicating a performance trade-off associated with deeper reasoning in simple questions. These trends are not present when context is not provided to the models, as shown in Supplementary Figure S13-A).
Figure 6: Analysis of the impact of the number of hops on models’ performance in Context Provided setup. panel A illustrates the relationship between output token usage (x-axis) and correctness rate (y-axis) for varying numbers of hops. Each point represents a model, while the colored areas indicate distribution estimates for each hop count. Panel B, depicts the distribution of correctness rates for non-reasoning models across questions with different numbers of hops, with the dots representing the median of each distribution.
In non-reasoning models, the output token count is not sensitive to the context or complexity of the question; thus, these models are evaluated solely based on answer correctness. Figure 6-B depicts the distribution of answer correctness rate across the evaluated reasoning models for different numbers of hops. Single-hop questions are answered at a higher correctness rate than multi-hop questions, a trend not observed as clearly in the context-free setup (Supplementary figure S13-B).
@@ 6 Conclusion
In this study, we developed a domain-specific multi-hop question-answering (QA) system and evaluated state-of-the-art large language models within the chemistry domain.Our findings reveal that these models struggle with in-domain multi-hop scientific questions, correctly answering fewer than half of the queries when context is unavailable. Although reasoning fine-tuned models show marginally improved performance, they still face significant challenges. The incorporation of context leads to substantial enhancements, nearly doubling the performance of both reasoning and non-reasoning models. However, even with context, no model, including those fine-tuned for reasoning, achieved a perfect score. Additionally, we contribute to the field by proposing an automated pipeline that integrates advanced named entity recognition with knowledge graph construction to generate intricate multi-hop reasoning tasks, which were utilized for the benchmark. Notably, this potentially domain-agnostic framework can be adapted for various fields by adjusting the pipeline to suit the target domain – such as replacing chemistry-specific named entity recognition with appropriate alternatives– thereby laying a robust foundation for future research aimed at enhancing reasoning capabilities across diverse specialized domains.
@@
Like all research, our study has certain limitations. Our benchmarking was conducted in two setups: without context and with full context provided. However, real-world applications typically involve step-by-step context collection throughout the reasoning process, leading to the possibility of partial context retrieval. Non-reasoning models often rely on a single retrieval round prior to generating answers, which does not reflect this iterative nature. Implementing a retrieval-augmented generation (RAG) pipeline with multi-step retrieval could better align with practical scenarios, bridging the gap between oracle and no-context performance. Recent advancements in reasoning models have explored integrating generative models with RAG systems for multi-step retrieval [28, 38, 39, 40]. Future work should focus on developing and validating an industrial-grade, multi-step retrieval pipeline specifically for chemical texts, leveraging incremental context acquisition and structured reasoning to enhance accuracy in addressing complex chemical queries.
@@ Acknowledgements
The author(s) gratefully acknowledge the financial support for the research, authorship, and/or publication of this article provided by MITACS under funding number IT32409. We also thank Adam Wojciech Bartwiki for his leadership and project management expertise; Tobias Roth for his essential contributions to infrastructure; and Stephen Dokas for his invaluable recommendations in chemistry.
S7.1 Detailed Knowledge Graph Generation
In this section, we explain each step in our graph generation process, illustrating how unstructured chemical text is transformed into a structured representation suitable for downstream tasks.
S7.1.1 Text Preprocessing
Our preprocessing pipeline is designed to isolate the most informative portions of ChemRxiv articles and prepare them for entity and relation extraction. First, we retrieved all ChemRxiv articles whose licenses permit redistribution via the ChemRxiv API. Using regular expressions, we extracted the introduction section of each paper, as this typically focuses on concise, objective statements about chemical phenomena. From each introduction, we extracted the first 500 words; this limit balances coverage of key background information against the risk of including less relevant or overly general text. To avoid fragmenting paragraphs, we ensured that no paragraph was split across chunks. Finally, we segmented each introduction into contiguous chunks of up to 128 words, preserving natural sentence boundaries. This segmentation reduces the input length for subsequent processing steps, ensuring that each chunk remains within the token limits of our downstream models while retaining semantic coherence.
S7.1.2 Node Extraction
Once the text is segmented, we identify chemical entities and their interrelations. We apply a named entity recognition (NER) model based on the PubMedBERT architecture, fine-tuned on multiple chemical datasets, to detect candidate entities within each 128-word chunk. To further improve precision and eliminate spurious mentions, we refine the NER output using OpenAI’s gpt-4o, prompting it with a few shots to verify whether each detected span indeed corresponds to a chemically meaningful entity and to standardize entity labels (for instance, converting "MeOH" to "methanol"). Below, we showed the prompt for Entity verification.
You are a chemistry expert specializing in entity recognition. Your task is to validate and filter the extracted entities, ensuring they are chemically meaningful based on the provided text. Remove any irrelevant terms, including general descriptors, numerical values, reaction conditions, and vague terms.
Entities Extracted by NER:
{entities}
Text for Context:
{text}
Criteria for Valid Entities:
  • $\checkmark$
    Chemical compounds (e.g., "HCl", "Sodium hydroxide", "Ethanol", "Benzene")
  • $\checkmark$
    Chemical elements (e.g., "Carbon", "Oxygen", "Cesium")
  • $\checkmark$
    Specific catalysts, solvents, reagents (e.g., "Cs2CO3", "Toluene", "Palladium")
Remove the Following Types of Entities:
  • $\times$
    Generic terms (e.g., "Reaction", "Solvent", "Acid", "Base", "Solution")
  • $\times$
    Experimental conditions (e.g., "pH", "Temperature", "2 M", "Strong acid")
  • $\times$
    Measurement terms (e.g., "X-ray diffraction", "NMR")
  • $\times$
    General descriptors (e.g., "High concentration", "Low efficiency")
Output Format:
Return only a Python list of valid chemical entities, with no explanations, markdown, or extra formatting.
S7.1.3 Edge Extraction
After obtaining a vetted set of entities, we extract pairwise relations using the same gpt-4o instance. For each pair of entities co-occurring within a chunk, we prompt the model with a few shots to classify or generate the nature of their relationship, producing triplets of the form (entity_A, relation, entity_B). The result of this step is a set of entity nodes (chemical names) and directed edges (relation labels) extracted directly from the text, forming a raw triplet collection that reflects the functional associations present in the chemical literature. Below, we show the prompt for Edge Extraction.
You are an expert in chemical text analysis. Your task is to extract only chemically meaningful relationships between a given set of entities from the provided text.
Guidelines for Relation Extraction:
  1. 1.
    Entity Matching: Consider only the entities provided in the given set. If an entity appears in the text but has no meaningful chemical relationship with another entity in the set, ignore it.
  2. 2.
    Chemically Significant Relations Only: Extract relations that describe actual chemical interactions, transformations, or properties (e.g., "reacts with," "catalyzes," "dissolves in," "produces").
  3. 3.
    Factual Relations: Only extract factual relations. Avoid observations, opinions, and findings.
  4. 4.
    Tuple Format: Output extracted facts in the form of (entity1, relation, entity2).
  5. 5.
    Avoid Generic Relations: Exclude weak relations like "is," "are," "exists," "relates to." Focus on specific interactions.
Valid Relation Types (Examples):
  • $\checkmark$
    "reacts with"
  • $\checkmark$
    "catalyzes"
  • $\checkmark$
    "binds to"
  • $\checkmark$
    "dissolves in"
  • $\checkmark$
    "oxidizes"
  • $\checkmark$
    "inhibits"
  • $\checkmark$
    "precipitates with"
  • $\checkmark$
    "acts as a solvent for"
  • $\checkmark$
    "is synthesized from"
Avoid These Weak Relations: Exclude relations such as "is," "are," "has," "exists."
Entities Provided:
{entities}
Text:
{text}
Extract at most {max_facts} factual statements.
Output Format:
Provide the output as a Python list of tuples, containing only the extracted relationships without any code formatting, backticks, or markdown.
Example Output:
  • [ ("HCl", "dissolves in", "Water"), ("HCl", "reacts with", "Sodium hydroxide") ]
S7.1.4 Knowledge Enrichment
To enrich each node with descriptive metadata and resolve naming ambiguities, we retrieved supplemental information from two external resources: Wikipedia and PubChem. From Wikipedia, we extracted the introductory summary of each entity’s page, providing a concise description of its common usage, historical context, or primary function. From PubChem, we obtained the official compound name (Record Title) alongside additional names and identifiers drawn from the "Names and Identifiers" section. We retrieved a textual description from the "Record Description" heading, summarizing key properties or applications. Safety annotations, including hazard statements or pictograms, were collected from the "Chemical Safety" subsection. Structural information was captured in the form of the canonical SMILES string under "Computed Descriptors," and the molecular formula was fetched from the "Molecular Formula" field. Finally, we extracted computed physicochemical properties, such as molecular weight, topological polar surface area, and log P values, from the "Computed Properties" list within "Chemical and Physical Properties." All of these metadata fields were stored as additional text and added as external information for each of the nodes.
S7.1.5 Graph Generation
At this stage, we constructed the graph by forming triplets of the form (node, edge, node). Each node was linked to the original text segment from which it was extracted and, if available, to its corresponding external metadata. Likewise, each edge was associated with the specific text chunk that produced it. In this way, both nodes and edges maintain direct references to their source text, ensuring traceability throughout the knowledge graph.
S7.2 Detailed Question Generation
S7.2.1 Path Sampling from the Knowledge Graph
To generate multi-hop questions, we first sampled paths of varying lengths from the constructed knowledge graph using a randomized breadth-first search (BFS) path sampling algorithm. During path sampling, we enforced that each edge in a sampled path originates from a distinct source text, thereby encouraging the integration of information from multiple, separate context passages. Concretely, a path of length $K$ comprises $K+1$ entities and traces through $K$ different source texts extracted from the original ChemRxiv database. By requiring unique sources for each edge, we ensure that correct solutions must draw on evidence scattered across several documents rather than a single paragraph.
S7.2.2 One-Hop Question Formulation
Adopting a bottom-up approach, we generated individual one-hop questions for each edge along a sampled path. For every triplet $(\text{entity}_{1},\text{relation},\text{entity}_{2})$, we framed a question whose answer is $\text{entity}_{1}$ by asking, "Which entity holds the relation relation to $\text{entity}_{2}$?" If the initial phrasing lacked sufficient specificity or clarity, we invoked a language model (OpenAI’s o3-mini) to augment the prompt with metadata drawn from the original text segment, such as contextual phrases or qualifying details. This enrichment step ensures that each one-hop question remains clear, precise, and answerable from the associated source text alone. Below is the prompt for generating the one-hop questions.
You are given a text along with an entity and its relation to another entity.
Entity 1: {entity1}
Relation: {relation}
Entity 2: {entity2}
Text: {text}
Information about Entity1: {entity1_meta if entity1_meta else None}
Your task is to generate a factual question whose answer is Entity1.
The question should ask for the entity that has the specified relation to Entity2.
Do not mention the answer (which is Entity1) in the question.
Ensure that the question is factual and can be answered solely based on the given text and the information about Entity1.
Do not refer to sections such as "Abstract," "Table #1," "in the text," or "in the article."
If Entity1 and relation are not specific enough (i.e., multiple answers are possible), add descriptions from the text or from the information about Entity1 to make it specific so that Entity1 is the only answer.
Return a dictionary without any code formatting, backticks, or markdown, with keys "q" and "a".
S7.2.3 Multi-Hop Question Aggregation
Once all one-hop questions for a path were vetted, we combined them into a single multi-hop question by prompting OpenAI’s o3-mini model, with a few shots. The final aggregated question is constructed by starting with the sub-question corresponding to the last edge (i.e., the entity nearest to the "tail" of the path) and then chaining backward through each preceding entity until reaching $\text{entity}_{1}$ of the first relation. In other words, if the sampled path is
$(\text{entity}_{1}\xrightarrow{\text{relation}_{1}}\text{entity}_{2}),\;(\text{entity}_{2}\xrightarrow{\text{relation}_{2}}\text{entity}_{3}),\;\dots,\;(\text{entity}_{K}\xrightarrow{\text{relation}_{K}}\text{entity}_{K+1}),$
the aggregated question asks first about $\text{entity}_{K+1}$ (the final tail), then uses its answer as context for the penultimate question, and so on, so that the final answer corresponds to $\text{entity}_{1}$. This reverse-chaining structure ensures that the multi-hop prompt leads directly to the original target node while preserving logical flow. Below, we show the prompt for generating the multi-hop question.
You are given multiple factual questions and their answers that are logically connected.
Your task is to chain them into a single, coherent multi-hop question that requires multiple reasoning steps.
Ensure that the (only) answer is the answer to the first question, and the question naturally follows from the facts given.
You have to start from the last generated question and build up a single multi-hop question so it aggregates them all and the answer is the answer to the first question.
None of the answers to any of the questions should be in the generated question.
Here is an example:
Example:
Q1: What is oxidized to form Carbon Dioxide?
A1: Methane
Q2: What is used in Photosynthesis?
A2: Carbon Dioxide
Q3: What produces Oxygen?
A3: Photosynthesis
Multi-hop question:
Q: What is oxidized to produce a substance that is used in a process that results in Oxygen?
A: Methane
Here are the generated questions and answers:
{formatted_qas}
Return a Python dictionary without any code formatting, backticks, or markdown, with keys "q" (multi-hop question) and "a" (final answer).
S7.2.4 Verification and Filtering
During verification, we first reviewed each one-hop question for clarity, chemical relevance, and direct answerability based on its corresponding source text. Below is the prompt for one-hop question evaluation.
You are a chemistry expert. Your task is to determine if the given question is a factual chemistry question, unambiguous (has only one answer), and answerable based on the provided context. A factual question must be based on actual chemical properties, reactions, or experimentally verified principles and must be strictly related to chemistry. An answerable question should be solvable based on the given context and must not be open-ended or have multiple correct answers. MAKE SURE THE QUESTION HAS ONLY ONE CORRECT ANSWER. There shouldn’t be any other entity except for the given answer that could be another answer.
### Question:
{question}
### Answer:
{answer}
### Context:
{context}
Please analyze the context and verify if the question is factual, unambiguous, and answerable. If the question is factual, has only one correct answer, is strictly related to chemistry, and can be answered based on the context, return "yes." Otherwise, return "no."
### Examples of Factual Chemistry Questions:
✓"What dissolves in water and evaporates at 0 °C?"
✓"What catalyst is used in the reaction between A and B?"
### Examples of Non-Factual or Ambiguous Chemistry Questions:
✗✗"What is the song of Nirvana that is a chemical entity?"
✗✗"What chemical entity and structural unit form the layered hydroxide structures with intercalated water ions used in battery materials and OER catalysis?" (M(OH)6 and $\alpha$-Ni(OH)2 are valid answers)
✗✗Questions that have multiple possible correct answers or are not strictly related to chemistry.
Next, the multi-hop question underwent an additional evaluation step to confirm that the logical chain, from the final sub-question back to the first, correctly guides a reader (or model) to $\text{entity}_{1}$. We employed an LLM-based verification process to assess factual accuracy, answerability given available context and metadata, and overall coherence among sub-questions. Feedback from domain experts was continuously incorporated into the prompt templates to refine verification accuracy. Any question, either one-hop or multi-hop, that was answered incorrectly by all evaluated models was excluded from the benchmark to minimize ambiguity and ensure high-quality, unambiguous reasoning tasks. Below is the prompt for evaluating the whole path.
You are a chemistry expert. Your task is to determine if the given question is a factual chemistry question and answerable based on the provided path.
### Path Information:
{path_text}
### Question:
{question}
### Answer:
{answer}
Please analyze the path and verify if the question is a factual chemistry question and can be answered based on the given path. A factual question must be based on actual chemical properties, reactions, or experimentally verified principles. An answerable question should be solvable based on the given path. If the question is factual and answerable, return "yes". If it contains speculation, opinions, or lacks verifiable chemical grounding, or it is not solvable, return "no".
### Examples of Factual Chemistry Questions:
✓"What dissolves in water?"
✓"What catalyst is used in the reaction between A and B?"
✓"Which compound undergoes oxidation in this reaction?"
✓"What product is formed when sodium reacts with chlorine?"
### Examples of Non-Factual Chemistry Questions:
✗✗"Why do some scientists think this reaction is inefficient?"
✗✗"What is the best solvent for this reaction?"
✗✗"Is this reaction useful in industry?"
✗✗"Do you think this compound is a good catalyst?"
Provide only "yes" or "no" as your response.
S7.2.5 Short-Answer Design and Rationale
To minimize the impact of writing style and summarization on accuracy evaluation, all questions were deliberately designed to elicit short, concise answers. Answering a multi-hop question requires decomposing it into its constitutive one-hop sub-questions, retrieving each intermediate answer, and then combining them to arrive at the final answer. Even when full context is available, a correct response cannot be produced if a model fails to infer and integrate multiple pieces of knowledge. By keeping answers brief and focusing on discrete factual steps, we ensure that performance evaluation reflects a model’s ability to conduct multi-hop inference rather than its capacity to paraphrase or generate lengthy explanations.
S7.3 Rejected Questions
There were a couple of reasons that we identified for rejecting of some questions from the benchmark. One was having multiple valid answers, and second was the answer was present in the question. Some examples are provided below.
Q1:
Context:
Researchers have developed an anode material based on NiCorGO (Nickel–Cobalt–reduced Graphene Oxide). In some variants, the NiCorGO is further decorated with palladium (Pd) nanoparticles to enhance catalytic performance.
Question:
Which component in the electrode structure functions as a catalyst at the anode when incorporating decorated NiCorGO?
Issue:
The question is declined due to ambiguity; two distinct answers are technically correct based on the variant of the material:
  • If the material is NiCorGO without Pd, the catalyst is Nickel (Ni).
  • If the material is Pd-decorated NiCorGO, the catalyst is Palladium (Pd).
Since the phrasing of the question does not clearly disambiguate which material variant is being used, it leads to multiple valid interpretations. Therefore, it cannot be accepted as a single-answer question.
Q2:
Context:
Hypervalent iodine compounds such as diaryliodonium salts are widely used as electrophilic arylation reagents. According to the source text, these salts are employed in both transition metal-catalyzed and metal-free arylation reactions. These reactions can be used to functionalize aromatic compounds, including halogen-substituted analogues like CIMPPC, by replacing hydrogen atoms.
Question:
Which type of arylation reaction that employs electrophilic arylation reagents utilizes diaryliodonium salts in hypervalent iodine chemistry?
Expected Answer:
Transition metal-catalyzed arylations
Reason for Rejection:
The question is declined due to the presence of multiple valid answers. The source explicitly states that diaryliodonium salts are used in:
  • Transition metal-catalyzed arylations, and
  • Metal-free arylations.
Both are equally valid interpretations of the question. Without further constraints or clarification, the question has more than one correct answer and does not meet the single-answer requirement.
To minimize ambiguity, we excluded questions that were answered incorrectly by all evaluated models from the benchmark. A subset of these questions was manually assessed, with most categorized as having multiple valid answers. Additionally, as explained in Section 3.2, an overall LLM verifier evaluated all questions for quality and answer inclusion. Any questions rejected by this verifier were also removed from the pool.
S7.4 Expert feedback
As summarized in Table S4, we report the average number of models that produced a correct answer with full context versus no context for each category. We also include the mean number of reasoning hops per question (2.46–2.60). Overall, the observations indicate that question complexity, measured by the number of hops, and the accuracy of LLM answers do not reliably predict question quality. All annotation guidelines, per-question ratings, and raw model outputs are available in our public repository.
Cat. Num. Que. Avg. Corr. (+ctx) Avg. Corr. (–ctx) Avg. len.
Good 26 (65%) 7.5 4.2 2.46
Ok 9 (22.5%) 7.55 3.44 2.55
Poor 5 (12.5%) 7.8 4.6 2.60
Table S4: The 40 high-confidence questions are grouped by expert-rated quality (Good, Ok, Poor). Num. Que. shows the count and percentage of questions in each category. Avg. Corr. (+ctx) / (–ctx) gives the mean number of models that answered correctly when provided with full supporting context versus no context, highlighting how contextual evidence boosts multi-hop inference. Avg. len. denotes the average number of reasoning hops required per question.
S7.5 Detailed Performance Based on Context Availability
@@
Figure S7 shows that model performance is strongly influenced by whether context is provided in the input. In particular, Claude 3.7 with extended thinking achieved a correctness rate of 84% with context, whereas o3-mini recorded the highest correctness rate (48%) when the context was absent. Note that o3-mini was primarily used to generate the questions, which may have introduced a slight bias, resulting in its minor improvement in correctness. Figures S8 and S9 illustrate the token usage and latency of the models, respectively.
@@
S7.6 Performance of models on Chemistry Subset of HotpotQA
Table S5 shows the details of each model’s performance, latency, and tokens used in both setups of context provided and not provided for the chemistry subset of HotpotQA [19] questions.
@@
Table S5: Summary of tested models’ performance on HotpotQA chemistry subset in terms of several evaluation metrics for both Contextual and Non-Contextual Setups
S7.7 A Multi-Hop QA Generation Example
@@
[Source 1*]: Carbonylation reactions constitute a potent tool to manufacture carboxylic acids and their derivatives both in industry and academic organic synthesis. In general, carbonylation requires the use of toxic carbon monoxide, which thus usually demands certified high-pressure reaction vessels. Therefore, developing non-gaseous CO surrogate for conducting safe and facile-operation carbonylation is an important and ongoing research topic. Among these established CO surrogates, formic acid is one kind of versatile atom.
[Source 2*]: The utilization of carbon dioxide as a C1 feedstock for the generation of industrially relevant chemicals is also an interesting approach. CO2 is an attractive renewable C1 source, which can lead to formic acid. Those approaches would not only reduce carbon dioxide emissions through carbon capture but also compensate sequestration costs by producing chemicals in global demand. Question:
@@
1) formic acid can be produced from carbon dioxide.
2) formic acid is the simplest carboxylic acid with antibacterial and preservative properties. 3) formic acid can act as a non-gaseous CO surrogate. 4) carbonylation reactions safely produce carboxylic acids under mild conditions using formic acid as a CO surrogate. Path (multi-hop chain of reasoning):
@@
* Source 1 and source 2 are coming from different documents.
@@
Context:
[Source 1*]: The most common way to functionalise two-dimensional materials such as graphene is through reactions occurring in solution .
[Source 2*]: Radiolytic shielding via graphene membranes has been demonstrated, highlighting the role of the two-dimensional allotrope graphene. [Source 3*]: A variety of chemisorptionbased solid sorbents exist, such as amines grafted onto porous solids like silica, cellulose, or metalorganic frameworks (MOFs). Membrane-based DAC captures CO2 by selectively allowing CO2 to permeate membranes while excluding other gases like nitrogen . [Source 4*]: Cr3(Cr4Cl)3(BTT)82 (BTT3 1,3,5-benzenetristetrazolate)12  MOF exhibits very high selectivity for O2 over nitrogen . While multiple factors can influence gas adsorption in MOFs, the origin of very high selectivity for O2 is perplexing because structurally similar metal ion MOFs are unselective. Question:
What is the metal–organic framework that exhibits very high oxygen selectivity over the major atmospheric diatomic gas, which is typically excluded by the selective barriers used for isolating carbon dioxide, and which act as radiolytic shields when formed from a two-dimensional carbon allotrope that is often functionalised in solution?
Answer: Cr3(Cr4Cl)3(BTT)82 (BTT3 1,3,5-benzenetristetrazolate)12 Sentence-level supporting facts:
1) Solution-phase chemistry is the standard route to functionalise graphene.
2) Graphene can form membranes that provide radiolytic shielding. 3) Membranes used for DAC selectively exclude nitrogen. 4) Cr3(Cr4Cl)3(BTT)82 MOF shows very high O2 selectivity over nitrogen. Path (multi-hop chain of reasoning):
solution
$\rightarrow$ graphene $\rightarrow$ membranes $\rightarrow$ nitrogen $\rightarrow$ Cr3(Cr4Cl)3(BTT)82
*Sources 1–4 are extracted from four different documents.
Figure S11: A 4-hop multi-document question–answer example.
S7.8 Impact of Context and Reasoning on output tokens count
Figure S12 illustrates the impact of context availability on the average number of output tokens generated by reasoning and non-reasoning models when answering questions. Non-reasoning models produce a similar number of tokens, as they do not engage in test-time reasoning. In contrast, for reasoning models, the number of tokens generated to answer questions decreases with the availability of context, indicating a potential requirement for less thinking when the context is available.
@@
S7.9 Performance Analysis Based on Number of Hops
Figure S13 illustrates how the number of hops affects performance in the absence of context. The data reveals that as the number of hops increases, the correctness rate declines while the number of output tokens rises. Additionally, both Figure LABEL:fig:hops_context and Figure S13 show a negative correlation between token count and correctness when only one hop is used. This may suggest that overanalyzing simpler questions could lead to errors in the final answer. We visualized the overall performance of all models in Figure S14 and  S15 to analyze how the correctness rate varies with the number of reasoning hops. In addition, token usage and latency metrics were separately depicted in Figures S16,  S17,  S18 and S19, respectively, to provide a more detailed view of the efficiency and resource consumption as the reasoning depth increases.
Figure S13: Analysis of the impact of the number of hops on the model’s performance in the absence of context. A) Scatter and KDE plot of the number of hops vs. correctness rate and output tokens. B) Distribution of the non-reasoning models’ performance for different numbers of hops.
The following figures illustrate the details of each model’s performance, latency, and tokens used for question clusters requiring a different number of hops to be answered.